OpenStudy (anonymous):

One number is 1 less than a second number. Twice the second number is 20 less than 4 times the first. Find the smaller of two numbers.

OpenStudy (dumbcow):

Let 1st number be A, 2nd num be B A = B-1 2B = 4A -20 --> 2B = 4(B-1) -20 2B = 4B -24 B = 12 --> A = 12-1 = 11 smaller num is 11

OpenStudy (campbell_st):

the 2 numbers are x and (x-1) then 2(x-1) = 4x -20

OpenStudy (campbell_st):

2x - 2 = 4x - 20 18 = 2x x = 9 then x-1 = 8

OpenStudy (anonymous):

Thanks